27. On December 31, 2020, the partnership of Abe, Bravo, Charlie and Delfin decided to liquidate with the following account balances. Cash NCA Liab. Bravo Charlie Abe Bravo Charlie Delfin Loan Loan (10%) (20%) (45%) (25%) 20,000 180,000 100,000 10,000 5,000 15,000 25,000 15,000 30,000 Assuming that Bravo and Charlie are limited partners and Abe, Bravo, Charlie and Delfin are solvent up to P10,000; P15,000; P10,000 and P5,000 respectively; determine the cash settlement given to Abe if the NCAS were sold for P100,000 (round to nearest peso).


1
Expert's answer
2021-12-17T10:25:17-0500

Solution:


Notes:

  • Loss on sale of NCA is allocated to all partners based on their profit and loss ratios.
  • Additional investment of Charlie is only P10,000 because he is solvent up to P10,000 only, as stated.
  • The remaining debit balance of Charlie's Capital after his loan and additional investment is absorbed by the remaining partners and allocated on the basis of their P/L ratio.

Abe - 10/55 "\times"6,000 = 1,091

Bravo - 20/55"\times" 6,000 = 2,182

Delfin - 25/55 "\times" 6,000 = 2,727